# A Framework of SO(3)-equivariant Non-linear Representation Learning and its Application to Electronic-Structure Hamiltonian Prediction

### v3

@@ B.1 Proof of Theorem 1
Thus, the product $\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}$ is:
$\mathbf{Q}(\mathbf{R})\cdot\mathbf{Q}(\mathbf{R})^{\dagger}=(\mathbf{D}^{l_{p}}(\mathbf{R})\cdot\mathbf{Q}\cdot\mathbf{D}^{l_{q}}(\mathbf{R})^{\dagger})\cdot(\mathbf{D}^{l_{q}}(\mathbf{R})\cdot\mathbf{Q}^{\dagger}\cdot\mathbf{D}^{l_{p}}(\mathbf{R})^{\dagger}).$
@@
B.2 Proof of Theorem 2
Proof.
The input feature $\mathbf{f}$ is SO(3)-equivariant, meaning that under an SO(3) rotation represented by $\mathbf{R}$, it transforms as follows:
@@ B.2 Proof of Theorem 2
First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar because the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations. Since applying a non-linear operation to an SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$. It formally holds that:
$z=z(\mathbf{R})$ (9)
Next, we apply the chain rule in Jacobian form. Considering $\mathbf{f}(\mathbf{R})$ is a column vector, to facilitate the application of the chain rule in vector form, we first transpose it into a row vector $\mathbf{f}(\mathbf{R})^{T}$, then differentiate:
@@
Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$. Since the representations of neural networks are generally real numbers, the Wigner-D matrix is also real unitary  [^3] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, because the output space may be complex while the internal representations of neural networks are typically real. Although the internal representations of neural network modules including $g_{\text{nonlin}}(\cdot)$ are in the real field, they are still able to predict targets in complex field with equivariance. Literature [8] has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts..
@@
Appendix C Information of Experimental Databases
@@
In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH series [6, 8], and the two databases from the QH9 series [19], listed in Table 5 and Table 6, respectively. Additionally, we visualize two types of challenging testing samples: samples with non-rigid deformation from thermal motions, as well as the bilayer samples with interlayer twists, which are shown in Fig. 4 and Fig. 5, respectively.
@@
Table 5: Statistic information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series [6, 8]. SOC: effects of Spin-Orbit Coupling. $m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset. $nt$: non-twisted samples. $t$: twisted samples.
@@
Table 6: Basic statistics of the two benchmark databases, QH9-Stable ($QS$) and QH9-Dynamic ($QD$), from the QH9 benchmark series [19]. For the $QS$ database, the ’ood’ strategy from the official settings is used to split the training, validation, and test sets, ensuring that the atom number of samples do not overlap across the three subsets. For the $QD$ database, the ’mol’ strategy from the official settings is employed to split the training, validation, and test sets, ensuring that there are no thermal motion samples from the same temporal trajectory across the three subsets. These split strategies aim to assess the regression model’s extrapolation capability with respect to the number of atoms as well as the temporal trajectories. $m$: number of samples in the current dataset. $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset.
@@
Appendix D Training Configurations
$\lambda$ in the training loss function is set as $0.3$ according to parameter selection on the validation sets. The initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$ and $BS$ databases are set as $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively, the training batch size for them is set as $1$ and the optimizer is selected as Adam; the initial learning rate for experiments on the $QS$ and $QD$ databases is set as $0.0005$, the training batch size is set as $16$, and the optimizer is selected as AdamW.
Appendix E Visualization of Block-level MAE Statistics
The MAE metrics (denoted as $MAE^{H}_{block}$) for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, across different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig. 6 and 7. Fig. 6 presents the results for monolayer structures, while Fig. 7 focuses on bilayer structures. From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, i.e., DeepH-E3, on most of the blocks of the Hamiltonian matrix, especially on those blocks where DeepH-E3 performs poorly.
Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the $MG$ and $MM$ databases.
Figure 7: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) structures of $BG$, $BB$, $BT$ and $BS$.
Appendix F Ablation Study
@@
• DeepH-E3 [8]: the baseline model.
• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include the gradient-based mechanism for encoding SO(3)-equivariant non-linear features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression.
• DeepH-E3+Grad: this setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term. Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to yield SO(3)-equivariant non-linear features based on the representations provided by the baseline encoders of DeepH-E3 for Hamiltonian prediction. As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, predicting only the Hamiltonian without joint supervision training through the trace label.
@@
Table 7: Ablation study MAE results (meV) on the $MG$ and $MM$ databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Table 8: Ablation study MAE results (meV) on $BG$, $BB$, $BT$and $BS$. The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Experimental results of the four setups are listed in Table 7 and 8. From the results of ablation terms, we can obtain a more fine-grained experimental analysis. By comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually. Moreover, their combination provides even better performance.

### v4

@@ B Proofs of Theorems
Proof of Theorem 2.
Under the given condition, the input feature $\mathbf{f}$ in direct-sum state is SO(3)-equivariant, meaning that under an SO(3) rotation represented by $\mathbf{R}$, it transforms as follows:
@@
First, according to group theory, $u=\text{CGDecomp}(\mathbf{f}\otimes\mathbf{f},0)$ is an SO(3)-invariant scalar as the degree-zero component from the Clebsch-Gordan decomposition is invariant under rotations. Since applying a non-linear operation to an SO(3)-invariant quantity does not change its invariance, $z=s_{\text{nonlin}}(u)$ is also SO(3)-invariant, independent to the specific form of $s_{\text{nonlin}}(\cdot)$. It formally holds that:
$z(\mathbf{R})=z$ (9)
Next, we apply the chain rule in Jacobian form. Considering $\mathbf{f}(\mathbf{R})$ is in the form of a column vector, to facilitate the application of the chain rule in vector form, we first transpose it into a row vector $\mathbf{f}(\mathbf{R})^{T}$, then differentiate:
@@
Here we utilize the property that $\mathbf{D}^{l}(\mathbf{R})^{-1}=\mathbf{D}^{l}(\mathbf{R})^{T}$  [^2] In Theorem 1 and Theorem 2, the Wigner-D matrices are in the complex and real fields, respectively, since the target quantity may be complex, whereas the internal representations of neural networks are typically in the real field. Nonetheless, neural network representations in the real field can still predict complex-valued targets with SO(3)-equivariance. Previous literature (Gong et al., 2023) has provided mechanisms for converting the network outputs in the real field into regression targets with real and imaginary parts.. Since the representations of neural networks are generally real numbers, the corresponding Wigner-D matrix is also real unitary.
@@
C Information of Experimental Databases
Figure 2: Illustration of a Hamiltonian matrix. Each Hamiltonian matrix contains multiple basic blocks, with the $\mathbf{Q}=\mathbf{Q}^{l_{p}\otimes l_{q}}$ defined in Section 3 corresponding to a basic block $\mathbf{H}^{l_{p}\otimes l_{q}}$ here.
@@ C Information of Experimental Databases
In this part, we provide detailed information about the experimental databases, including the statistical information of the six databases from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023) and the two databases from the QH9 benchmark series (Yu et al., 2023a), listed in Table 4 and Table 5, respectively. Additionally, we visualize two types of challenging testing samples: samples with non-rigid deformation from thermal motions, as well as the bilayer samples with interlayer twists, which are shown in Fig. 3 and Fig. 4, respectively.
D Implementation Details
The hardware environment for our experiments is a server cluster equipped with Nvidia RTX A6000 GPUs, each with 48 GiB of memory. Other experimental details may differ across the DeepH and QH9 benchmark series, which we will describe separately.
E Implementation Details on the DeepH Benchmark Series
The software environment used is Pytorch 2.0.1 for experiments on the six crystalline databases from the DeepH benchmark series. When combining the proposed TraceGrad method with the DeepH-E3 architecture, the implementation of DeepH-E3 is based on the project  [^3] https://github.com/Xiaoxun-Gong/DeepH-E3 provided by Gong et al. (2023), keeping the architecture and hyperparameter configurations consistent with their setup. In our framework, we use the same number of encoding modules $K$ as DeepH-E3, which is set to $3$. In each encoding module, we apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature, enabling non-linear expressiveness. We set the number of channels for the rotationally SO(3)-invariant feature $\textbf{u}^{(k)}$ $(1\leq k\leq 3)$ to $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$; the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $25$ for $MG$ and $BG$, $49$ for $MM$, and $196$ for $BB$, $BT$, and $BS$. It is worth noting that, while $s_{nonlin}(\cdot)$ can be implemented as any differentiable neural network module, we here implement it as a simple fully-connected module. This decision is made to avoid adding significant computational burden to the whole network. Meanwhile, as DeepH-E3 already incorporates complex graph network mechanisms for information aggregation and message-passing, there is no need for $s_{nonlin}(\cdot)$ to be overly complex. Its role is focusing on to filling in the gaps left by the existing equivariant mechanisms in DeepH-E3: to introduce a non-linear mapping mechanism that maintains equivariance, thereby activating and unleashing the expressive power of the overall network architecture through non-linearity. The SO(3)-equivariant decoder we adopt is the same as that of DeepH-E3; The SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $3$ $(K)$ times of the dimensionality of $\textbf{z}^{k}$, e.g., $75$ for $MG$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix. Since each basic block of the Hamiltonian matrix can compute a trace, the total number of trace variables corresponds to the number of basic blocks. Regarding the error metric in the loss function Eq. 7, for the first term, we follow DeepH-E3 to use MSE (Mean Squared Error); for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on crystalline structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the six crystalline databases and the searched value is $0.3$. Other hyper-parameters and configurations are the same as DeepH-E3 (Gong et al., 2023): the initial learning rates for experiments on the $MG$, $MM$, $BG$, $BB$, $BT$, and $BS$ databases are set to $0.003$, $0.005$, $0.003$, $0.005$, $0.004$, and $0.005$, respectively; the training batch size is set as $1$; the optimizer is chosen as Adam; the scheduler is configured as a slippery slope scheduler.
F Implementation Details on the QH9 Benchmark Series
The software environment used is Pytorch 1.11.0 for experiments on the two molecular databases from the QH9 benchmark series. When combining the proposed TraceGrad method with the QHNet architecture, the implementation of QHNet is based on the project  [^4] https://github.com/divelab/AIRS provided by Yu et al. (2023b), keeping the architecture and hyperparameter configurations consistent with their setup. In our framework, we use the same number of encoding modules as QHNet: $5$ node feature encoding modules and $2$ edge feature encoding modules. We opt to apply the $g_{nonlin}(\cdot)$ module proposed in Section 4 and 5 to each SO(3)-equivariant edge feature. We set the number of channels for $\textbf{u}^{(k)}$ $(1\leq k\leq 2)$ as $1024$. The neural network module $s_{nonlin}(\cdot)$ within $g_{nonlin}(\cdot)$ is implemented as a three-layers fully-connected module: the input size is set to $1024$, consistent with $\textbf{u}^{(k)}$, the hidden layer size is also $1024$, with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism, and the output layer size (i.e., the dimensionality of $\textbf{z}^{(k)}$) is set to be equal to the number of basic blocks for a Hamiltonian matrix, which is $36$ for $QS$ and $QD$ databases. The SO(3)-equivariant decoder we adopt is the same as that of QHNet; the SO(3)-invariant decoder we adopt is a four-layers fully-connected module: the input size is $2$ $(K)$ times of the dimensionality of $\textbf{z}^{(k)}$, e.g., $72$ for $QS$ and $QD$; the hidden layers have $1024$ neurons with SiLU as the non-linear activation function and LayerNorm as the normalization mechanism; the size of the output layer is the number of basic blocks for a Hamiltonian matrix. Regarding the error metric in the loss function Eq. 7, for the first term, we follow QHNet to use a combination of MSE and MAE; for the second term, we choose between MSE and MAE based on performance on the validation sets, ultimately selecting MAE. $\lambda$ in the training loss function is set according to parameter selection on the validation sets, searching from $\{0.1,0.2,...,1.0\}$. Here, we aim to obtain a more general parameter setting for $\lambda$ on molecular structures, and thus we determined $\lambda$ based on the overall performance on the validation sets of the two molecular databases and the searched value is $0.2$. Other hyper-parameters and configurations are the same as QHNet: the initial learning rates for all experiments are set as $5\times 10^{-4}$, the training batch size is set as $32$, the optimizer is selected as AdamW, and a learning rate scheduler is implemented. To be more specific, the scheduler gradually increases the learning rate from 00 to a maximum value of $5\times 10^{-4}$ over the first $1,000$ warm-up steps. Subsequently, the scheduler linearly reduces the learning rate, ensuring it reaches $1\times 10^{-7}$ by the final step.
G Visualization of Block-level MAE Statistics
As shown in Fig. 2, each Hamiltonian matrix consists of numerous basic block, with each basic block representing the direct product of two degrees. Here, we follow Yin et al. (2024) to measure the MAE performance of deep models on each basic block, denoted as $MAE^{H}_{block}$. The values of $MAE^{H}_{block}$ for the two setups, i.e., DeepH-E3 and DeepH-E3+TraceGrad, on different blocks of the Hamiltonian matrix for six databases from the DeepH benchmark series are illustrated in Fig. 5 and 6. Fig. 5 presents the results for monolayer structures, while Fig. 6 focuses on bilayer structures. From these figures, it can be observed that our method, TraceGrad, brings significant accuracy improvements over the baseline method, DeepH-E3, across the vast majority of blocks of the Hamiltonian matrices, particularly on blocks where DeepH-E3 struggles with lower accuracy.
@@ G Visualization of Block-level MAE Statistics
Table 4: Statistical information of the six benchmark databases, i.e., Monolayer Graphene ($MG$), Monolayer MoS2 ($MM$), Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), Bilayer Bi2Se3 ($BS$), from the DeepH benchmark series (Li et al., 2022; Gong et al., 2023). SOC: effects of Spin-Orbit Coupling. $m$: number of samples in the current dataset; $a_{max}$: maximum number of atoms from a unit cell in the current dataset. $a_{min}$: minimum number of atoms from a unit cell in the current dataset. $nt$: non-twisted samples. $t$: twisted samples.
@@
Table 5: Statistical information of the two benchmark databases, QH9-stable ($QS$) and QH9-dynamic ($QD$), from the QH9 benchmark series (Yu et al., 2023a). The $QS$ database is split using the ’ood’ strategy, while the $QD$ database is split using the ’mol’ strategy. $m$: number of samples in the current dataset. $a_{max}$: maximum number of atoms for a sample in the current dataset. $a_{min}$: minimum number of atoms for a sample in the current dataset.
@@
Figure 5: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases.
Figure 6: Visualization of $MAE^{H}_{block}$ on each basic block of the Hamiltonian matrices for the non-twisted (marked with superscripts $nt$) and twisted (marked with superscripts $t$) testing subsets of Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$).
H Ablation Study
@@
• DeepH-E3 (Gong et al., 2023): the baseline model.
• DeepH-E3+Trace: this experimental setup, an ablation term, only implements half part of our method. Specifically, it extends the architecture of DeepH-E3 by adding our SO(3)-invariant encoding and decoding branches and using the trace quantity $\mathbf{T}=\text{tr}(\mathbf{Q}\cdot\mathbf{Q}^{\dagger})=\text{tr}(\mathbf{H}^{l_{p}\otimes l_{q}}\cdot(\mathbf{H}^{l_{p}\otimes l_{q}})^{\dagger})$ to train them. As for ablation study, this setup does not include the gradient-based mechanism delivering non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features; instead, it directly uses the SO(3)-equivariant features outputted by DeepH-E3 for Hamiltonian regression. In this configuration, the SO(3)-invariant branches only contribute indirectly during the training phase by backpropagating the supervision signals from the trace quantity to the earlier layers.
• DeepH-E3+Grad: this setup is also an ablation term and implements the other half part of our method in contrast to the previous ablation term. Specifically, it incorporates our SO(3)-invariant encoder branch as well as the gradient-induced operator to deliver non-linear expressiveness from SO(3)-invariant features to encode SO(3)-equivariant features. As for ablation study, this setup continues to use the single-task training pipeline of DeepH-E3, supervised only with the Hamiltonian label without joint supervised training through the trace of Hamiltonian.
@@
Table 6: Ablation study MAE results (meV) on the Monolayer Graphene ($MG$) and Monolayer MoS2 ($MM$) databases. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Table 7: Ablation study MAE results (meV) on the Bilayer Graphene ($BG$), Bilayer Bismuthene ($BB$), Bilayer Bi2Te3 ($BT$), and Bilayer Bi2Se3 ($BS$) databases. The superscripts $nt$ and $t$ respectively denote the non-twisted and twisted subsets. $\downarrow$ means lower values of the metrics correspond to better accuracy.
@@
Experimental results of the four setups are listed in Table 6 and 7. Table 6 presents the results for monolayer structures, while Table 7 focuses on bilayer structures. From the results of ablation terms, we can obtain a more fine-grained experimental analysis. By comparing among the results of DeepH-E3, DeepH-E3+Trace, DeepH-E3+Grad, and DeepH-E3+TraceGrad, we can conclude that the two core mechanisms of our method, i.e., the SO(3)-invariant trace supervision mechanism (Trace) at the label level as well as the gradient-based induction mechanism (Grad) at the representation layer, can contribute to the performance individually. Moreover, their combination provides even better performance. This is because, on one hand, with the gradient-based induction mechanism as a bridge, the non-linear expressiveness of SO(3)-invariant features learned from the trace label can be transformed into the SO(3)-equivariant representations during inference; on the other hand, with trace label, the SO(3)-invariant network branch has a strong supervisory signal, enabling it to learn the intrinsic symmetry and complexity of the regression targets, enhancing the quality of SO(3)-invariant features and ultimately benefits the encoding of SO(3)-equivariant features. The value of such complementarity has been fully demonstrated in the experimental results.
I Discussion on Time Costs
When the atomic system is small and any two atoms are within the cutoff radius of each other such that the Hamiltonian is defined, the number of Hamiltonian matrices is approximately proportional to $N^{2}$, where $N$ is the number of atoms in the system. Yet, as the atomic system grows, the system exhibits a dual nature of locality and non-locality. Locality refers to the fact that, when $N$ becomes large, the number of atoms within the cutoff radius of each atom remains roughly constant, and thus the growth in the number of Hamiltonian matrices asymptotically scales as $O(N)$. On the other hand, the non-locality arises from the fact that, although Hamiltonian matrices are defined only for pairs of atoms within each other’s cutoff radius, atoms outside the cutoff radius may still exert long-range effects on the Hamiltonian matrices within the cutoff radius
The baseline models we select, whether DeepH-E3 (Gong et al., 2023) or QHNet (Yu et al., 2023b), are SO(3)-equivariant graph neural network models with efficient information aggregation and message-passing mechanisms, which cleverly balance the locality and non-locality of the system. As a result, the computational complexity asymptotically scales as $O(N)$ as $N$ grows to large. This is a significant efficiency advantage compared to the $O(N^{3})$ computational complexity of traditional DFT methods, enabling efficient and low-cost prediction of electronic structures for large-scale atomic systems. Our method improves the Hamiltonian prediction accuracy of DeepH-E3 and QHNet while maintaining this advantage of asymptotic complexity. The neural module $g_{nonlin}(\cdot)$ we introduce efficiently updates each edge feature $\textbf{f}^{(k)}$ $(1\leq k\leq K)$ from the baseline models to obtain the corresponding non-linear SO(3)-equivariant feature $\textbf{v}^{(k)}$, ensuring that the computational complexity remains asymptotically $O(N)$, with good potential to handle large atomic systems at acceptable computational costs.
We test the inference time costs of the DeepH-E3 and DeepH-E3+TraceGrad models on an Nvidia RTX A6000 GPU, using the testing sets from two databases, namely $MG$ and $MM$, as examples; and test the inference time costs of the QHNet and QHNet+TraceGrad models on $QS$ and $QD$, on the same device. The batch size is set to 1 for all tests on inference time. Each sample is inferred 50 times, and the average time per inference is recorded. When using DeepH-E3, the average inference time per sample (s/sample) for $MG$ and $MM$ are as follows: $0.247$ and $0.256$, respectively; with DeepH-E3+TraceGrad, the average inference times are $0.264$ and $0.274$, correspondingly. When using QHNet, the average inference time per sample (s/sample) for $QS$ and $QD$ are as follows: $0.233$ and $0.174$, respectively; with QHNet+TraceGrad, the average inference times are $0.248$ and $0.187$, correspondingly. As we can see, the increases in computational costs due to the TraceGrad module are slight relative to that of the baseline models. Considering the significant accuracy improvements brought by the TraceGrad method as reported in Section 6.2, we believe that this minor increase in computational cost is acceptable and worthwhile in practical applications.
J Empirical Study on Combining our Method with Approximately SO(3)-equivariance Framework
While non-strict SO(3)-equivariance, which may limit the depth of theoretical exploration, is not the main focus of this study aiming at bridging rigorous SO(3)-equivariance with the non-linear expressive capabilities of neural networks, considering that they remain of interest in a few numerical computation applications where precision is highlighted over strict equivariance, we also conduct empirical study combining our method with approximately SO(3)-equivariant techniques. Taking the Monolayer MoS2 ($MM$) database as a case study, we evaluate the performance of combining our trace supervision and gradient induction method (TraceGrad) with the an approximately equivariant approach HarmoSE (Yin et al., 2024). We here take HarmoSE as the backbone encoder, and yields features by TraceGrad to enrich its representations. The experimental results in Table 8 and Fig. 7 demonstrate that TraceGrad significantly enhances the accuracy of HarmoSE, surpassing DeepH-2 and achieving SOTA results. Both DeepH-2 and HarmoSE sacrificed strict SO(3)-equivariance to fully release the expressive capabilities of graph Transformers, aiming for the ultimate in prediction accuracy. Despite this, our method still manages to significantly exceed their accuracy, further confirming the superiority and generality of our method in learning expressive representations of physical systems.
Figure 7: Comparison on the $MAE^{H}_{block}$ metric for HarmoSE and HarmoSE+TraceGrad on the $MM$ database.
Table 8: MAE results (meV) for DeepH-2, HarmoSE, and HarmoSE+TraceGrad on the $MM$ database. $\downarrow$ means lower values of the metrics correspond to better accuracy. The results of the compared methods are taken from the corresponding literature (Wang et al., 2024b; Yin et al., 2024), where the empty items are due to the data not being provided in the original paper.
Methods $MM$
$MAE$ ()($\downarrow$)( ↓ )
$MAE^{H}_{all}$ $MAE^{H}_{cha\_s}$ $MAE^{H}_{cha\_b}$
DeepH-2 (Wang et al., 2024b) 0.21 - -
HarmoSE (Yin et al., 2024) 0.233 0.293 0.406
HarmoSE+TraceGrad 0.178 0.228 0.296
K Future Work
First, from a theoretical and methodological perspective, we can generalize SO(3) group symmetry to a broader range and more complex groups in deep learning research. This will enable deep learning frameworks to accommodate richer mathematical structures and physical quantities. Second, from the application perspective, the current methodology extends beyond predicting electronic-structure Hamiltonians and is capable of generally predicting multiple physical quantities and properties that are equivariant. Furthermore, as our approach offers a promising general solution to the critical dilemma between equivariance and non-linear expressiveness for the deep learning paradigm, the methodology is not confined to physical research alone and may find applications in areas such as robotics, autonomous vehicles, and motion tracking systems. For instance, in the vision tasks of autonomous vehicles (Wang et al., 2023), where there is a high demand for robustness against coordinate transformations of cameras, the mainstream approach has been to achieve approximate SO(3)-equivariance through data augmentation. However, this does not guarantee absolute reliability and safety. Our work holds great potential to construct deep models with high generalization performance under the premise of strict SO(3)-equivariance, potentially revolutionizing techniques in such tasks. Looking forward to future endeavors, the foundational and pivotal role of this work is clear.